OpenStudy (anonymous):

a coin collection made up of pennies, nickles, and quarters contains total of 29 coins. the number of quarters is 8 less than the number of pennies. the total face value of the coins is $1.77. how many of each denomination are there?

OpenStudy (anonymous):

1) 29 coins: 29 = q + n + p 2) eight fewer quarters: q = p - 8 3) total value 177 cents: 177 = 25q + 10n + p

OpenStudy (anonymous):

is like that?